# Mutual Information Free Topological Generalization Bounds via Stability

### v1

@@ 2.2 Algorithmic stability
Let $\mathcal{A}(S,U)\in\mathbb{R}^{d}$ be a randomized algorithm where $S\in\mathcal{Z}^{n}$ is a dataset and $U$ represent the algorithm randomness. We say that $\mathcal{A}$ satisfies uniform parameter stability with coefficient $\beta>0$ if for any neighboring datasets $S\simeq S^{\prime}$, we have:
$\mathbb{E}_{U}\left[\left\|\mathcal{A}(S,U)-\mathcal{A}(S^{\prime},U)\right\|\right]\leq\beta.$
If $\ell(w,z)$ is $L$-Lipschitz continuous in $w$, this implies the classical algorithmic stability [HRS16]:
@@
Unfortunately, in the context of worst case generalization bounds (i.e., upper bounds on $G_{S}(\mathcal{W}_{S})$), this assumption is not of much use as one needs to specify the stability of the entire set $\mathcal{W}_{S}$ and not just one parameter inside it. To address this issue, a first step was taken by [FGK+19], who introduced the following notion of hypothesis set stability.
Definition 2.2 ([FGK+19]).
A family $(\mathcal{W}_{S})_{S\in\mathcal{Z}^{n}}$ of data-dependent hypothesis sets is said to be $\beta$-uniformly stable if for any $S\simeq S^{\prime}$ in $\mathcal{Z}^{n}$ and any $w\in\mathcal{W}_{S}$, if there exists $w^{\prime}\in\mathcal{W}_{S^{\prime}}$ such that:
@@
Note that the classical algorithmic stability (as described in [HRS16] for instance) is a particular case of Definition 2.2. The correctness of this assumption in several settings was studied in [FGK+19], such as strongly convex optimization. In Section 3, we will extend this condition to noisy algorithms.
In the context of topological generalization bounds, a different stability assumption was already used in [DDS23] to obtain generalization bounds with slightly simpler information-theoretic terms, but without removing the need for such terms. The main technical novelty of our study is to exploit stability to prove IT-free topological generalization bounds, as we present in the next section.
@@ 3 Main Theoretical Results
We now present our main theoretical contributions, which are expected topological generalization bounds on the generalization error $G_{S}(\mathcal{W}_{S})$ defined in Equation 3. Our core technical contribution is based on a variation of the hypothesis set stability of [FGK+19] (see Definition 2.2), which we present in Section 3.1. First, we specify some notation. Let $\mathrm{CL}(\mathbb{R}^{d})$ denote the closed subsets of $\mathbb{R}^{d}$ and $\Omega$ denote an arbitrary probability space, we see a learning algorithm as a mapping:
$\mathcal{A}:\bigcup_{n=1}^{\infty}\mathcal{Z}^{n}\times\Omega\longrightarrow\mathrm{CL}(\mathbb{R}^{d}),$ (7)
which we denote $\mathcal{A}(S,U)=\mathcal{W}_{S,U}$, where $S$ is the dataset and $U$ is an independent random variable representing the algorithm randomness. In practice, $\mathcal{W}_{S,U}$ is the trajectory of the optimizer in the parameter space, as specified in Section 1. This framework has been adopted by previous studies [HSKM22, DDS23]. Note that the dependence on $U$ will be omitted when the context is clear.
3.1 Random trajectory stability
The main drawback of Definition 2.2 is that it does not explicitly account for the algorithm randomness (e.g., batch indices), as opposed to the classical algorithmic stability [BE02, HRS16].
In order to address this issue, we propose a variant of the assumption of [FGK+19], accounting for the presence of randomness. The main difficulty that arises here is to give a meaning to the expressions such as $w\in\mathcal{W}_{S}$, when $\mathcal{W}_{S}$ is a random set. To do so, we define the notion of data-dependent selection [Mol17].
@@ 3.1 Random trajectory stability
A data-dependent selection of $\mathcal{W}_{S,U}$ is a deterministic mapping $\omega:\mathrm{CL}(\mathbb{R}^{d})\times\mathcal{Z}^{n}\to\mathbb{R}^{d}$ such that $\omega(\mathcal{W}_{S,U},S)\in\mathcal{W}_{S,U}$ almost surely with respect to the joint distribution of $(S,\mathcal{W}_{S,U})$.
We are mostly interested in the following “argmax selection”. In the rest of the paper, given $S_{1},S_{2}\in\mathcal{Z}^{n}$, we assume the existence of a random variable $\omega(\mathcal{W}_{S_{1}},S_{2})$ such that, almost surely:
$\omega(\mathcal{W}_{S_{1}},S_{2})\in\operatorname*{arg\,max}_{w\in\mathcal{W}_{S_{1}}}G_{S_{2}}(w).$ (8)
Note that the existence of such a random variable is ensured under mild conditions [Mol17]. Equipped with this definition, we now introduce our stability assumption.
Assumption 3.1 (Trajectory stability).
Algorithm $\mathcal{A}$ is $\beta_{n}$-trajectory-stable if for any $J\in\mathbb{N}^{\star}$ and any data-dependent selection $\omega$ of $\mathcal{W}_{S,U}$, there exists a mapping $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that:
  • almost surely, for any $S$ and $w$, $\omega^{\prime}(W_{S},w)\in\mathcal{W}_{S}$
@@
In the absence of algorithmic randomness, that is, when each $U$ is constant, 3.1 is a particular case of Definition 2.2. Stability assumptions are usually enounced for neighboring datasets, here we present a variant with datasets differing by $J$ elements. These two variants are equivalent and we chose this presentation to simplify subsequent proofs and notations. The presence of the mapping $\omega^{\prime}$ might be intriguing: $\omega^{\prime}$ corresponds to the parameter denoted $w^{\prime}$ in Definition 2.2 and is denoted this way to make explicit its dependence on the different random variables.
The main advantage of 3.1 is to take the expectation over the algorithm randomness. A natural question arises as to whether this assumption can be satisfied in practice. We answer this question by the following lemma, which shows that 3.1 is implied by Definition 2.1.
@@
This result shows that Definition 2.1 implies trajectory stability, in the sense of 3.1. While uniform argument stability differs from the classical notion of algorithmic stability, it can be noted that many stability-based studies rely on a Lipschitz assumption and prove uniform argument stability to derive algorithmic stability parameters [HRS16], as noted by [BFGT20]. Therefore, we see that the conditions of Lemma 3.2 can be satisfied in numerous cases, such as optimization of convex, smooth, and Lipschitz continuous functions [HRS16]. Under these conditions, it is obtained that the stability parameter is of order $\beta_{k}=\mathcal{O}(k/n)$, hence, yielding random trajectory stability with a parameter of order $\mathcal{O}(T^{2}/n)$, in the worst case. In the appendix, we present a corollary to Lemma 3.2 showing that projected SGD on a non-convex problem is $\mathcal{O}(1/n)$-trajectory-stable.
This finding is key to our approach. Indeed, combined with the topological generalization bounds as we do in the next section, a stability parameter of order $\mathcal{O}(1/n)$ will yield a complete bound of the form $\mathcal{O}(n^{-1/3}\sqrt{\mathbf{C}(\mathcal{W}_{S})})$, where $\mathbf{C}(\mathcal{W}_{S})$ is a topological complexity term.
3.2 Topological generalization bounds for trajectory-stable algorithms
We now leverage our random trajectory stability assumption to prove new topological generalization bounds without IT terms. We assume that the trajectories $\mathcal{W}_{S}$ are almost surely finite throughout, i.e., they correspond to the trajectory of the optimizer for a finite number of iterations.
As in previous work [DVDS24, ADS+24], our proofs are centered around the notion of Rademacher complexity [BM02] (formally defined in the appendix). One of our key technical contributions is deriving a Rademacher complexity bound for trajectory-stable algorithms. For the sake of simplicity, we postpone the technical details to the appendix and present our main results hereafter.
Our new topological generalization bounds rely on regularity conditions of the loss function (namely, Lipschitz continuity and boundedness), which we now present.
Assumption 3.3.
The loss function $\ell\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}$ is $L$-Lipschitz continuous with respect to its first argument, i.e., for all $w,w^{\prime}\in\mathbb{R}^{d}$ and any $z\in\mathcal{Z}$, $|\ell(w,z)-\ell(w^{\prime},z)|\leq L\|w-w^{\prime}\|$.
Such Lipschitz assumptions were already used in previous topology-based studies [SSDE20, BLGS21, ALRS23]. Subsequent works [DDS23] mitigated these assumptions, albeit through the incorporation of intricate IT terms, which we avoid in our stability-based framework. As is typical in this literature [ADS+24, DVDS24, BLGS21], we also assume the boundedness of the loss function.
@@ 3.2 Topological generalization bounds for trajectory-stable algorithms
We remark that this assumption holds; for instance, for the 00-$1$ loss. As demonstrated in [DDS23, ADS+24], the proposed theory remains experimentally valid even for unbounded losses. Our experimental results support extending this observation to our setting.
Equipped with these assumptions, we can now present our new stability-based topological generalization bounds. We present two results, depending on the topological complexity of interest. The first result is a topological bound based on $\alpha$-weighted lifetime sums.
@@
Suppose that Assumptions 3.1, 3.3, and 3.4 hold and define $K_{n,\alpha}:=2(2L\sqrt{n}/B)^{\alpha}$. Without loss of generality [^5] Not making this assumption only yields changes in the absolute constants of the bound., assume that $\beta_{n}^{-2/3}$ is an integer divisor of $n$. For any $\alpha\in(0,1]$, we have:
$\mathbb{E}\bigg{[}\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\bigg{]}\leq\beta_{n}^{1/3}\left(2+2B+2B\mathbb{E}\left[\sqrt{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}_{\left\|\cdot\right\|}(\mathcal{W}_{S,U}))}\right]\right).$
The detailed proof is postponed to the appendix. By building on the results on positive magnitude of [ADS+24], this leads to the following positive magnitude-based generalization bound for trajectory-stable algorithms.
@@
Suppose that Assumptions 3.1, 3.3, and 3.4 hold and define $s(\lambda):=\lambda L\beta_{n}^{-1/3}/B$ Without loss of generality, assume that $\beta_{n}^{-2/3}$ is an integer divisor of $n$. For any $\lambda>0$ we have:
$\mathbb{E}\bigg{[}\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\bigg{]}\leq\beta_{n}^{1/3}\left(2+\lambda B+\frac{2B}{\lambda}\mathbb{E}\left[\log\mathbf{PMag}_{\left\|\cdot\right\|}\left(s(\lambda)\cdot\mathcal{W}_{S,U}\right)\right]\right).$
This result suggests that a theoretically justified choice of the magnitude scale would be $s(\lambda)\approx\beta_{n}^{-1/3}$, which amounts to $\Theta(n^{1/3})$ in the event that $\beta_{n}=\mathcal{O}(1/n)$. This is an interesting observation, as the results in [ADS+24] suggested an optimal scale parameter of order $s=\Theta(\sqrt{n})$.
Note moreover that it was also observed by that small values of $s(\lambda)$ can give satisfactory correlation with the generalization error, indicating that better choices of $\lambda$ might be available in certain cases. We leave the investigation of this matter to future studies.
In Theorems 3.5 and 3.6, we focus on two topological complexities: the weighted lifetime sums and the positive magnitude. This choice is motivated by the empirical results of [ADS+24], who demonstrated that they have better correlation with the generalization error compared to previously introduced dimension-based complexities [SSDE20]. Nevertheless, our approach can trivially recover bounds based on other complexities, such as the persistent homology dimension [BLGS21].
Our bounds show that for trajectory-stable algorithms, the topological complexities (i.e., lifetime sums and positive magnitude) have an important impact on the worst-case generalization error. As opposed to existing trajectory-dependent topological bounds [SSDE20, BLGS21, DDS23, ADS+24], our bounds do not feature any IT terms. In addition to replacing IT terms with a stability assumption, the stability parameter $\beta_{n}$ acts as a multiplying constant in the bound, hence increasing the impact of the topological complexity on the bound. Compared to the IT terms, $\beta_{n}$ is easily interpretable and can be expected to decrease with $n$ at a polynomial rate, as discussed above.
Despite being the first mutual information-free topological bounds, Theorems 3.5 and 3.6 have three main disadvantages: (i) we only provide expected bounds, (ii) we obtain a slightly worse convergence rate in $n$ and (iii) our framework only allows for Euclidean-based topological complexity, and we typically do not cover the case of the data-dependent pseudometric introduced by [DDS23]. Regarding (ii), our results should be compared with [DDS23, Theorem 3.8], which employed a stability-like assumption to obtain smaller IT terms than existing works at the time. They obtained a convergence rate that is similar to Theorems 3.5 and 3.6 but our results do not feature any IT term.
Stability parameters are expected to be affected by the sample size, hence, these results suggest to study the correlation between generalization error and topological complexity for several values of $n$, which forms a part of our empirical study.

### v2

@@ 2 Technical Background
Let $\mathcal{A}:(S,U)\to\mathbb{R}^{d}$ be a randomized algorithm. $\mathcal{A}$ satisfies $\beta$-uniform argument stability if for any neighboring datasets $S\simeq S^{\prime}$, we have $\mathbb{E}_{U}\left[\left\|\mathcal{A}(S,U)-\mathcal{A}(S^{\prime},U)\right\|\right]\leq\beta$.
If $\ell(w,z)$ is $L$-Lipschitz continuous in $w$, this implies the algorithmic stability (Hardt et al., 2016):
@@
In the context of worst-case generalization bounds, where $\mathcal{W}_{S,U}$ is a set (i.e., not a singleton), one needs to specify the stability of the entire set $\mathcal{W}_{S,U}$. To address this issue, a first step was taken by (Foster et al., 2019), who introduced the following notion of hypothesis set stability.
Definition 2.2.
A family $(\mathcal{W}_{S})_{S\in\mathcal{Z}^{n}}\in(\mathbb{R}^{d})^{\mathcal{Z}^{n}}$ of data-dependent hypothesis sets is $\beta$-uniformly stable when for any $S\simeq S^{\prime}$ in $\mathcal{Z}^{n}$ and $w\in\mathcal{W}_{S}$, there exists $w^{\prime}\in\mathcal{W}_{S^{\prime}}$ such that:
@@
Although this assumption has been analyzed in settings such as strongly convex optimization (Foster et al., 2019), the notion of hypothesis set stability (Definition 2.2) does not account for the presence of algorithmic randomness $U$ and therefore cannot be applied to the data-dependent random sets defined by Equation˜3. In Section 3, we thus extend Definition˜2.2 to data-dependent random sets.
@@ 3 Main Theoretical Results
In this section, we present our main theoretical contributions, which consist in deriving expected worst-case generalization bounds through a new concept of random set stability, described in Section˜3.1. After discussing its validity, we present a key technical lemma and recover classical generalization bounds in the case of Examples˜1.3 and 1.4. We then apply our framework to improve several data-dependent worst-case generalization bounds, including intrinsic dimension bounds (Simsekli et al., 2020; Birdal et al., 2021) and topological bounds (Andreeva et al., 2024).
3.1 Random set stability
As explained above, the main drawback of Definition˜2.2 is that it does not explicitly account for the algorithm randomness (e.g., batch indices), while algorithmic randomness is known to be paramount for single-iterates stability bounds (Hardt et al., 2016; Zhu et al., 2023).
To address this issue, we propose an improvement of the assumption of Foster et al. (2019), accounting for the presence of randomness. The main difficulty that arises here is to give a meaning to the expressions such as “for all $w\in\mathcal{W}_{S,U}$”, when $\mathcal{W}_{S,U}$ is a random set. To do so, we define the notion of data-dependent selection (Molchanov, 2017).
@@ 3.1 Random set stability
A data-dependent selection of $\mathcal{W}_{S,U}$ is a deterministic mapping $\omega:\mathrm{CL}(\mathbb{R}^{d})\times\mathcal{Z}^{n}\to\mathbb{R}^{d}$ such that $\omega(\mathcal{W}_{S,U},S^{\prime})\in\mathcal{W}_{S,U}$, almost surely. In particular, we assume the existence of a random variable $\omega_{0}(\mathcal{W}_{S,U},S^{\prime})$ such that, almost surely, $\omega_{0}(\mathcal{W}_{S,U},S^{\prime})\in\operatorname*{arg\,max}_{w\in\mathcal{W}_{S,U}}G_{S^{\prime}}(w)$, where we recall that $G_{S^{\prime}}(w):=\mathcal{R}(w)-\widehat{\mathcal{R}}_{S^{\prime}}(w)$ for all $S^{\prime}\in\mathcal{Z}^{n}$.
Note that the existence of such a random variable is ensured under mild measure-theoretic conditions (Molchanov, 2017). Equipped with this definition, we now introduce our stability assumption.
Assumption 3.1 (Random set stability).
Algorithm $\mathcal{A}$ is $\beta_{n}$-random set stable if for any $J\in\mathbb{N}^{\star}$ and any data-dependent selection $\omega$ of $\mathcal{W}_{S,U}$, there exists a map $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that:
  • For any $S$, $U$ and $w\in\mathbb{R}^{d}$, $\omega^{\prime}(\mathcal{W}_{S,U},w)\in\mathcal{W}_{S,U}$ .
@@
In the absence of algorithmic randomness (i.e., when $U$ is constant), ˜3.1 is a particular case of Definition˜2.2. Stability assumptions are usually enounced for neighboring datasets; here, we present a variant with datasets differing by $J$ elements. These two variants are equivalent, and we choose this presentation to simplify subsequent proofs and notations. The presence of the mapping $\omega^{\prime}$ might be intriguing: $\omega^{\prime}$ corresponds to the parameter denoted $w^{\prime}$ in Definition˜2.2 and is denoted this way to make its dependence on the different random variables explicit.
Applicability of random set stability
The question naturally arises as to whether this assumption can be met and how it fits within the existing stability conditions. We answer this question with the following lemma, which shows that ˜3.1 is implied by Definition˜2.1.
@@
In the context of Example˜1.1, this result shows that Definition˜2.1 implies random set stability. It can be noted that many stability-based studies rely on a Lipschitz assumption and prove uniform argument stability to derive algorithmic stability parameters (Hardt et al., 2016), as noted by Bassily et al. (2020). Therefore, we see that the conditions of Lemma 3.2 can be satisfied in numerous cases, such as optimization of convex, smooth, and Lipschitz continuous functions (Hardt et al., 2016). Under these conditions, it is obtained that the stability parameter is of order $\delta_{k}=\mathcal{O}(k/n)$, hence, yielding random set stability with a parameter of order $\mathcal{O}(T^{2}/n)$, in the worst case.
Projected SGD
Fix $T\in\mathbb{N}$. Let us consider the projected SGD recursion on a loss function $\ell$, which can be non-convex, i.e., given a dataset $S=(z_{1},\dots,z_{n})\in\mathcal{Z}^{n}$, consider the following algorithm:
$w_{k+1}=\Pi_{R}\left[w_{k}-\eta_{k}\nabla\ell(w_{k},z_{i_{k+1}})\right],\quad\mathcal{A}_{\mathrm{SGD}}(S,U)=\mathcal{W}_{S,U}:=\{w_{1},\dots,w_{T}\}$ (7)
where $i_{k}\sim\mathcal{U}(\{1,\dots,n\})$ are random indices, $U:=\{i_{k}\}_{k\geq 1}$, and $\Pi_{R}$ is the projection on the ball $\bar{B}_{R}(0)$. The next corollary uses Lemma˜3.2 to establish the random set stability of SGD for Lipschitz and smooth losses, and is an adaptation of (Hardt et al., 2016, Theorem 3.12) to our setting.
Corollary 3.3.
Suppose that, for all $z$, $\ell(\cdot,z)$ is $L$-Lipschitz, $\nabla\ell(\cdot,z)$ is $G$-Lipschitz, and that $\eta_{k}\leq c/k$ for some $c<1/G$. Then, $\mathcal{A}_{\mathrm{SGD}}$ is random-set-stable with parameter
$\textstyle\beta_{n}=\frac{4LR}{n-1}\left(\frac{L}{GR}\right)^{1/Gc+1}\sum_{1\leq k\leq T}k^{\frac{Gc}{Gc+1}}.$
3.2 Worst-case generalization bound for data-dependent stable random sets
We now leverage ˜3.1 to prove worst-case generalization bounds over random sets. The next lemma shows that the expected worst-case generalization error can be bounded in terms of the random set stability parameter and a Rademacher complexity term (Bartlett and Mendelson, 2002).
Lemma 3.4.
Let $S\sim\mu_{z}^{\otimes n}$. Let $J,K\in\mathbb{N}^{\star}$ be such that $n=JK$, and $\tilde{S}_{J}\sim\mu_{z}^{\otimes J}$ be independent of $S$ and $U$. Assume that algorithm $\mathcal{A}$ satisfies ˜3.1 with parameter $\beta_{n}>0$, then:
$\textstyle\mathbb{E}\big[\sup_{w\in\mathcal{W}_{S,U}}\big(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\big)\big]\leq 2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right]+2J\beta_{n},$ (8)
where $\mathbf{Rad}_{S}(\mathcal{W})$ denotes the Rademacher complexity, whose definition is recalled in Definition˜A.1.
Most existing data-dependent worst-case generalization bounds rely on an information-theoretic approach (Simsekli et al., 2020; Hodgkinson et al., 2022; Andreeva et al., 2024). In particular, this theorem can be seen as an analog of (Dupuis et al., 2024, Theorem 6), which provided a bound in terms of Rademacher complexity and information-theoretic terms. Instead of information-theoretic terms, our framework makes use of the random set stability parameter $\beta_{n}$.
In Lemma˜3.4, the number $J$ is a free parameter. Before discussing the main applications of our methods, we show that a certain choice of $J$ recovers: (1) classical algorithmic stability bounds ($J=1$) and (2) worst-case bounds over fixed hypothesis sets ($J=n$). Therefore, the parameter $J$ can be seen as interpolating between these two settings, while yielding new data-dependent worst-case generalization bounds for intermediary values of $J$.
Corollary 3.5 (Recovering stability bounds).
Consider an algorithm $\mathcal{A}$ as in Example˜1.3 and assume that $\mathcal{A}$ satisfies ˜3.1 with parameter $\beta_{n}$. Then, $\mathcal{A}$ is algorithmically stable in the sense of Equation˜6. By setting $J=1$ in Lemma˜3.4, we obtain:
$\mathbb{E}\big[\mathcal{R}(\mathcal{A}(S,U))-\widehat{\mathcal{R}}_{S}(\mathcal{A}(S,U))\big]\leq 2\beta_{n}.$
Up to the absolute constant $2$, this recovers classical algorithmic stability-based expected generalization bounds (Bousquet and Elisseeff, 2002; Bassily et al., 2020).
Corollary 3.6 (Recovering bounds on fixed hypothesis sets).
Consider the setting of Example˜1.4. Then, the algorithm satisfies ˜3.1 with parameter $\beta_{n}=0$. By using $J=n$ we obtain:
$\textstyle\mathbb{E}\big[\sup_{w\in\mathcal{W}}\big(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\big)\big]\leq 2\mathbb{E}\left[\mathbf{Rad}_{S}(\mathcal{W})\right],$
where we observe that $\tilde{S}_{n}\sim\mu_{z}^{\otimes n}$ and $S\sim\mu_{z}^{\otimes n}$ have the same distribution.
We observe that when $\beta_{n}=\mathcal{O}\left(n^{-1}\right)$, Corollary˜3.5 recovers the classical convergence rate of Hardt et al. (2016). Corollary˜3.6 shows that our framework tightly recovers the classical Rademacher complexity bounds (for fixed hypothesis set $\mathcal{W}$) (Bartlett and Mendelson, 2002), as well as all known consequences of these bounds, such as VC dimension bounds (Vapnik and Vapnik, 1998) and covering number bounds (Shalev-Shwartz and Ben-David, 2014). In particluar, Corollary˜3.6 shows that in the data-independent case we obtain the usual worst-case convergence rate of $\mathcal{O}\left(n^{-1/2}\right)$.
4 Applications
In this section, we apply our random set stability framework to improve recently proposed data-dependent worst-case generalization bounds in terms of fractal dimensions and topological complexities (Andreeva et al., 2024). We assume that a learning algorithm $\mathcal{A}(S,U)=:\mathcal{W}_{S,U}$ is given, in the sense of Equation˜3. One may think of $\mathcal{W}_{S,U}$ as a learning trajectory as in Example˜1.1.
Assumptions
Before presenting our main applications, we introduce some structural assumptions. We first note that the Rademacher complexity (RC) term appearing in Lemma˜3.4 is computed using a sample $\tilde{S}_{J}\sim\mu_{z}^{\otimes J}$ independent of the random set $\mathcal{W}_{S,U}$. Despite this, the set over which the RC term is computed is the data-dependent random set $\mathcal{W}_{S,U}$, while the independent sample $\tilde{S}_{J}$ is only used to evaluate the loss inside the RC. In order to exploit this property, we assume that the loss has a “local" Lipschitz regularity, in the following sense.
Assumption 4.1 (Lipschitz on random sets).
There exists a measurable map $(S,U)\mapsto L_{S,U}$ such that for all $S$ and all $U$, for all $w,w^{\prime}\in\mathcal{W}_{S,U}$, for all $z\in\mathcal{Z}$, $|\ell(w,z)-\ell(w^{\prime},z)|\leq L_{S,U}\|w-w^{\prime}\|$.
Lipschitz continuity assumptions are commonly used in the context data-dependent worst-case bounds (Simsekli et al., 2020; Birdal et al., 2021; Andreeva et al., 2023). However, our framework crucially only requires the Lipschitz continuity of $\ell(\cdot,z)$ to hold on each set $\mathcal{W}_{S,U}$ individually, with a Lipschitz constant $L_{S,U}$ that depends both on the dataset $S\in\mathcal{Z}^{n}$ and the noise $U$. Note however that this local Lipschitz continuity of $\ell(\cdot,z)$ is still required to be uniform in $z\in\mathcal{Z}$. For instance, in the case of Example˜1.1, where $\mathcal{W}_{S,U}$ is finite, then ˜4.1 is satisfied as soon as $\mathcal{Z}$ is finite (or compact if $z\mapsto\nabla\ell(w,z)$ is continuous for a topology on $\mathcal{Z}$).
As is typical in this literature (Andreeva et al., 2024; Dupuis et al., 2024; Birdal et al., 2021; Bartlett and Mendelson, 2002), we also assume the boundedness of the loss function.
@@ 4 Applications
4.1 Intrinsic dimensions and topological bounds
Recent studies (Simsekli et al., 2020; Dupuis et al., 2023; Andreeva et al., 2023; Dupuis et al., 2024; Birdal et al., 2021; Andreeva et al., 2024) have established upper-bounds on Equation˜4 and drawn links between the topological structure of $\mathcal{W}_{S,U}$ and the generalization error. These bounds are of the form presented in Equation˜5. As mentioned in Section˜1, one of the main drawbacks of these bounds is that they contain intractable and poorly understood mutual information terms. One of our main contributions is to show that, under random set stability, we are able to obtain data-dependent worst-case generalization bounds featuring many of the most popular complexity measures, without the need of the intractable mutual information terms (denoted IT in Equation˜5).
We first present an adaptation to our setting of the intrinsic dimension-based bounds of (Simsekli et al., 2020), without any mutual information term.
@@ 4.1 Intrinsic dimensions and topological bounds
Suppose that Assumptions 4.2, 3.1, and 4.1 hold, and that $\mathcal{W}_{S,U}$ is a.s. of finite diameter. Without loss of generality, assume that $\beta_{n}^{-2/3}$ is an integer divisor of $n$. There exists $\delta_{n}>0$ such that for all $\delta<\delta_{n}$
$\mathbb{E}\bigg[\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right)\bigg]\leq 2\mathbb{E}\bigg[\frac{B}{n}+\delta L_{S,U}+\beta_{n}^{1/3}\left(1+B\sqrt{4\overline{\dim}_{\mathrm{B}}(\mathcal{W}_{S,U})\log\frac{1}{\delta}}\right)\bigg],$
where $\overline{\dim}_{\mathrm{B}}(\mathcal{W}_{S,U})$ is the upper box-counting dimension (Falconer, 2013), which we define in the appendix in Definition˜A.4. We refer to the Section˜B.4 for a discussion of the parameter $\delta_{n}$.
In many cases, the upper box-counting dimension of $\mathcal{W}_{S,U}$ is much smaller than the ambient dimension $d$ (Falconer, 2013). Under mild assumptions, the upper box-counting dimension can be replaced by other notions like the persistent homology dimension (Birdal et al., 2021) and the magnitude dimension (Andreeva et al., 2023), which have provided promising empirical results.
However, it was argued by Andreeva et al. (2024) that such intrinsic dimensions are not well-suited to the practically used stochastic optimizers due to their discrete nature, as in Example˜1.1. To overcome this issue, these authors proposed to use two new notions of topological complexity, called the $\alpha$-weighted lifetime sums (denoted $\mathbf{E}^{\alpha}(\mathcal{W}_{S,U})$) and the positive magnitude (denoted $\mathbf{PMag}(s\cdot\mathcal{W}_{S,U})$), and significantly improved the empirical results. Intuitively, the $\alpha$-weighted lifetime can be understood as tracking the evolution of ‘connected components’ of a set across different scales and summing their contributions. On the other hand, the magnitude may be thought of as capturing the effective number of distinct points in a finite space at different scale Leinster (2013). We defer a precise definition of these notions to Definition˜A.6 and A.9.
In the next theorem, we provide IT free versions of the topological generalization bounds of Andreeva et al. (2024) under our assumption of random set stability (˜3.1).
@@
Suppose that Assumptions 4.2, 3.1, and 4.1 hold. We further assume that the set $\mathcal{W}_{S,U}$ is almost surely finite. Without loss of generality, assume that $\beta_{n}^{-2/3}$ is an integer divisor of $n$. Then, for any $\alpha\in(0,1]$, we have
$\mathbb{E}\bigg[\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\bigg]\leq\beta_{n}^{1/3}\left(2+2B+2B\mathbb{E}\left[\sqrt{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}(\mathcal{W}_{S,U}))}\right]\right),$
where $K_{n,\alpha}:=2(2L_{S,U}\sqrt{n}/B)^{\alpha}$. Moreover, for any $\lambda>0$, let $s(\lambda):=\lambda L_{S,U}\beta_{n}^{-1/3}/B$. We have
$\mathbb{E}\bigg[\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\bigg]\leq\beta_{n}^{1/3}\left(2+\lambda B+\frac{2B}{\lambda}\mathbb{E}\left[\log\mathbf{PMag}\left(s(\lambda)\cdot\mathcal{W}_{S,U}\right)\right]\right).$
This result suggests that a theoretically justified choice of the magnitude scale would be $s(\lambda)\approx\beta_{n}^{-1/3}$, which amounts to $\Theta(n^{1/3})$ in the event that $\beta_{n}=\mathcal{O}(1/n)$. As a reminder, in the classical setting Corollary˜3.6 we recover the standard convergence rate of $\mathcal{O}(n^{-1/2})$.
Unlike trajectory-dependent topological bounds (Simsekli et al., 2020; Birdal et al., 2021; Dupuis et al., 2023; Andreeva et al., 2024), our bounds do not involve IT terms, which can be unbounded. While our bounds result in a slower convergence rate, this represents a deliberate trade-off to maintain boundedness. In addition to replacing IT terms with a stability assumption, the stability parameter $\beta_{n}$ acts as a multiplying constant in the bound, hence increasing the impact of the topological complexity on the bound. Compared to the IT terms, $\beta_{n}$ is easily interpretable and can be expected to decrease with $n$ at a polynomial rate, as discussed above.